THE AREA OF A RECTANGLE IS 60 square METERS. FIND THE DIMENSIONS IF THE LENGTH IS 4 METERS LONGER THAN THE WIDTH"
Answer by TimothyLamb(4379)

a = Lw = 60
L = w + 4
---
Lw = 60
(w + 4)w = 60
ww + 4w - 60 = 0
---
the above quadratic equation is in standard form, with a=1, b=4 and c=-60
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 4 -60
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
w = 6
w = -10
---
the negative root doesn't fit the problem statement, so use the positive root:
---
answer:
w = 6 meters
L = 10 meters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php